OpenStudy (anonymous):

Did I answer this question properly? Find G(0) given g(x)= 4-8x. Substitute x in g(x) with 0. 0 = 4 - 8x Move 8x over. 8x=4 Divide 4 by 8. x=4/8 Simplify x=1/2 If I did this question incorrectly could you please tell me where I went wrong.

OpenStudy (akashdeepdeb):

We have to find G(0) So you don't substitute in place of G(x) You SUBSTITUTE x!! :D So it should be G(0) = 4 - 8(0) = 4 - 0 = 4!! So G(0) should be 4 !! :D Good try btw! :)

OpenStudy (ddcamp):

Almost, but you substituted 0 for g(x) instead of x: g(0) = 4 - 8(0) = 4-0 = 4

OpenStudy (anonymous):

Thank you

OpenStudy (akashdeepdeb):

:)